# Neural Networks for Fast Optimisation in Model Predictive Control: A Review

### v2

@@ 5 Neural Network Based Optimisation for Nonlinear Model Predictive Control
Parisini and Zoppoli (1995) 1a
First paper to propose the idea of training an NN to imitate MPC via pure regression.
None
6 3 2 30 30 Sim
Ortega and Camacho (1996) 1c
Train a shallow NN to imitate MPC using RMSE loss. Inputs to the network include past inputs and sensor readings that are not fed to MPC.
None
5 3 2 7 6 Robot
Cavagnari et al. (1999) 1e
Train shallow NN with sigmoid activation to imitate LMPC, generate training data from LQR solution as well as MPC.
None
4 4 1 10 10 Robot
Lucia et al. (2020) 1a
Train $tanh$ DNN with MSE loss to imitate MPC, then implement the approximate controller using FPGA for much faster execution.
None
2 2 2 10 10 500,000x HIL
Chan et al. (2021) 1b
Train DNN with MSE loss to imitate MPC, then add steady state error correction term to loss function and retrain to generate offset free DNN.
None
2 1 1 Sim
Vaupel et al. (2019) 1c
Train DNN to imitate MPC using MSE loss, use measurable disturbance variables as network inputs.
None
7 7 1 10 8 10x Sim
@@
Nurbayeva et al. (2022) 1c
MPC motion planner approximated by DNN via pure regression, explores input dimensionality reduction with autoencoders.
None
6 6 6 10 10 Robot
Löhr et al. (2019) 1a
Train DNN to imitate MPC with binary control inputs using RMSE loss function.
None
3 3 3 24 24 12.12x Sim
Karg and Lucia (2018) 1c
Train DNN to imitate MPC with binary control inputs. Include trajectories of disturbances over horizon as part of the training data.
None
4 4 4 24 24 Sim
Lovelett et al. (2020) 1c
Use manifold learning to map augmented state space to optimal control inputs. Then train DNN to retrieve optimal control inputs from manifold space.
None
2 1 1 20 20 36,000,000x Sim
Åkesson and Toivonen (2006) 1d
Train a shallow $tanh$ NN with an MPC loss function on data corresponding to closed loop trajectories. Use result to replace conventional optimiser.
None
4 2 2 60 25 Sim
Drgoňa et al. (2022) 1d
Create and train DNN plant based on state space architecture. Train DNN controller with loss function inspired on MPC cost.
None
1 1 32 32 Robot
Li et al. (2022) 2b
Formulate training data generation and DNN approximation training as a single optimisation problem.
Probabilistic constraint satisfaction.
4 2 2 20 20 1,000x Sim
Parisini et al. (1998) 2c
First formalised framework for generating shallow NN approximations of NMPC. Authors propose methods to compute NN size and terminal NMPC costs that make NN feasible.
Bound output error and stability to region near origin.
2 2 1 100 100 Sim
Pin et al. (2010) 2c
Further investigation of approximation properties of the framework proposed in Parisini et al. (1998).
Practical stability and input constraint satisfaction.
2 2 1 8 8 Sim
Pin et al. (2013) 2c
Extend  Pin et al. (2010) with requirements that approximations have to meet to guarantee stability and constraint satisfaction.
Practical stability and constraint satisfaction on inputs and states.
2 2 1 8 8 Sim
@@
Ahmed and Al-Dajani (1998) 1d
Create a first order differentiable, MPC-inspired loss function and use it train a DNN that imitates MPC. Derivatives of loss are computed using block partial derivatives.
Input constraint satisfaction via saturation and stability at equilibrium point.
2 1 1 300 300 Sim
Löhr et al. (2020) 3a
Train DNN to predict first value of a binary control input and thereby enable conversion of MIQP into QP.
Constraint satisfaction and input feasibility via solution of QP at every iteration.
3 3 5 48 48 3x Sim
Vaupel et al. (2020) 3a
Train DNN with MSE loss to imitate MPC then use it to hot start full NMPC solution or alongside a post processing QP that corrects the outputs of the network and makes them constraint compliant.
All MPC guarantees when DNN is used to hot start solver. Good practical constraint satisfaction when post processing QP is used.
6 5 3 7 5 24x Sim
Bonzanini et al. (2020) 3b
Train a DNN to imitate NMPC, project outputs of DNN onto RAI which in turn will ensure that states stay within RCI. The projection operation can be precomputed as the solution of an mpQP.
Input and state constraint satisfaction.
2 2 2 100 100 25 - 90x Robot
Wang et al. (2022) 1f
Integrate Lagrange duality into DNN loss function to guarantee constraint satisfaction of the primal problem, train the DNN approximation through imitation learning on the modified loss function.
Input and state constraint satisfaction.
3 3 2 50 50 6 - 72x Sim
@@ 5.1 Methods without guarantees
Parisini and Zoppoli Parisini and Zoppoli (1995) were the first to propose the idea of training an NN to imitate the outputs of an NMPC controller. Their framework consisted of fitting a shallow NN with $tanh$ activation to training data generated by repeatedly applying a stabilizing NMPC controller to initial states sampled randomly from a bound set. The inputs to the network were the current system states and the outputs were the control actions calculated by the NMPC controller for the first step in the control horizon, an MSE loss function and a gradient algorithm were used to solve the regression problem. The effectiveness of the approach was demonstrated via simulation on two example problems one of which included a MIMO plant of medium size as reported in Table 2. Ortega et al. Ortega and Camacho (1996) extended these results by applying the same method to a path tracking NMPC controller used as the navigation algorithm of a mobile robot. In this case the shallow NN was provided with additional inputs such as a parametric representation of the reference trajectory, distance sensor readings for obstacle detection, and most notably, the control actions at the previous time step. Ortega et al. Ortega and Camacho (1996) validated their approximate controller with experiments on the real robot, thus showing, for the first time, that NNBO can indeed facilitate the deployment of NMPC controllers that would otherwise be unable to execute in real time. Cavagnari et al. Cavagnari et al. (1999) further proved this point with their experimental results on a seesaw mechanism. Using the same random sampling and MSE loss framework, they replaced a stabilizing NMPC controller with an NN approximation capable of running in real time. Moreover, they introduced a small novelty by training the NN with data generated from two different controllers operating in different regimes. For samples near the equilibrium position, they used training data generated from an LQR controller with a linearised plant, for points outside this region, training data was generated using the NMPC controller. The result was an approximate NN controller that could seamlessly transition between two control behaviours for better stability. The three aforementioned studies were conducted in the late 90s. However, following their publication, the simple imitation learning approach received little attention for almost two decades.
@@
Interest in the simple imitation learning approach reignited around 2018, with many researchers proposing modifications and extensions of the method. Lucia et al. Lucia et al. (2020) showed that the $tanh$ DNN approximation of an NMPC controller could be implemented in Field Programmable Gate Array (FPGA) hardware to achieve much faster execution of the approximate controller and thus enable the control of systems with ultra fast dynamics. The effectiveness of this approach was demonstrated on a resonant power converter application including Hardware In the Loop (HIL) testing. From these tests, the authors Lucia et al. (2020) reported a speed up factor of approximately half a million times for their most optimised FPGA implementation compared to the original NMPC controller.
@@
Chan et al. Chan et al. (2021) proposed an extension of the basic framework to eliminate steady state setpoint tracking error due to approximation error. In their framework, a DNN approximation of an NMPC controller is first generated by solving the usual regression problem with an MSE loss function. Then, the approximation and original controller are applied to setpoint tracking trajectories to generate a new data set. This dataset is used to retrain the network using a modified loss function that includes an approximation error term. Retraining on this modified loss function effectively minimises the approximation error which in turn minimises, or effectively removes, setpoint tracking error. Chan et al Chan et al. (2021) also implemented their approximate controller on FPGA hardware and demonstrated the effectiveness of their approach via simulation for a small problem. Alternatively, Vaupel et al. Vaupel et al. (2019) showed that acceptable setpoint tracking performance can also be achieved by including disturbance measurements in the inputs of the neural network approximation and training with the pure regression setup on an MSE loss function. For applications where disturbances can be measured, this approach is significantly simpler compared to the method proposed by Chan et al. Chan et al. (2021).
@@
For complex systems with difficult to parameterise inputs, Nurbayeva et al. Nurbayeva et al. (2022) proposed the use of autoencoders for input dimensionality reduction. In their work, an NMPC motion planner for human-robot interaction tasks was replaced with a DNN approximation and successfully tested in simulation and on the real robot. Autoencoders were used to parameterise human motion inputs to the approximate controller, robot state inputs were fed directly. While the approach may be worth considering for some applications, the authors themselves point out that the resulting controller is not generalisable to motion scenarios beyond the specific interaction for which it was trained. The controller is also unable to cope with significant deviations from the interaction it was trained for.
@@
Löhr et al. Löhr et al. (2019) and Karg and Lucia Karg and Lucia (2018) showed that the basic imitation learning setup is also applicable to NMPC controllers with binary variables where the underlying optimisation problem is of the mixed-integer type. Löhr et al. Löhr et al. (2019) trained their approximation using an RMSE-based loss and the inputs to their network were the current states, references and disturbances of the system. Karg and Lucia Karg and Lucia (2018) considered the same inputs but used the entire predicted disturbance trajectory over the prediction horizon instead of only the latest measurement. In both studies the effectiveness of the approach was demonstrated via simulation on heating system control problems with MIMO plants of similar size. The independence and success of these studies suggests that NNBO is a viable method for accelerating NMPC problems with binary control variables, but there is room for in-robot validation studies as well as studies that explore the more general problem of handling integer variables instead of only binary variables.
@@
In a significantly different approach, Lovelett et al. Lovelett et al. (2020) noted that the mapping function between states and optimal control actions (e.g. the function that the NN in previous approaches is trying to learn) can be complicated and difficult to learn because the augmented state space, which consists of plant states and controller references, is often a poor parametrisation of the control law. Motivated by this observation, they proposed the use of diffusion maps, a type of manifold learning algorithm, to learn a potentially lower dimensional parametrisation of the mapping function. If such a parametrisation exists, it is a simpler map between manifold space and control actions. However, using this map to obtain control actions given the augmented state is not straightforward. The authors addressed this challenge with two NNs. The first NN was used to learn the mapping between augmented state space and manifold space, and the second network learnt the mapping between manifold space and optimal control actions. In theory these regression problems are easier than directly approximating the original controller. The authors demonstrated the effectiveness of the approach via a simulation case study on a nonlinear system with a small plant. Reports on execution time improvements indicated that the neural approximation was 36,000,000 times faster than the original NMPC.
@@
So far, we have discussed methods that make use of a simple imitation learning setup where error information is used to solve a regression problem via gradient descent, hereafter we present other alternatives. Åkesson and Toivonen Åkesson and Toivonen (2006) proposed an approach wherein the error-based loss function is replaced with an MPC-like loss function. This new loss function encourages the network to learn the dynamics of the controller instead of over-fitting it for a bound set of inputs. In their study the authors formulated this method for trajectory tracking NMPC. The network inputs are those typically available to an NPMC controller such as current estimated states, measured outputs, reference outputs, and previous control actions. The network outputs are the control command increments to be applied at the current control interval. In this case training consists of minimising the MPC-like loss function over a large set of trajectories on which the network is applied in a closed loop setup. The minimisation task can be formulated as a nonlinear least squares problem and solved using a gradient algorithm. The main advantage of this method is that the size of the resulting optimisation problem is independent of the horizon length used in the MPC loss function. This means that the NN approximation can be trained with longer horizons, surpassing what would be practically feasible for deployment in the exact NMPC controller.
@@
Drgoňa et al. Drgoňa et al. (2022) built further on the idea of departing from error-based loss functions in their Differentiable Predictive Control (DPC) framework where both the optimiser and plant model of the original controller are replaced with DNNs with architectures and loss functions inspired by MPC. More specifically, the plant model is approximated using a neural state space architecture which can learn the dynamics of nonlinear systems from recorded data. The main advantage of this approach is that the model directly learns the disturbed and uncertain dynamics of the system without additional effort. The trained plant can then be used to train a policy network using an MPC-inspired loss function and a dataset that emulates closed loop control. Crucially, the policy and plant networks can be combined into a single model that is continuously differentiable making it easy to train with any modern training algorithm such as ADAM Kingma and Ba (2014). Again, the advantage of combining the policy and data-derived plant networks into a single model is that the resulting policy inherently compensates for disturbances and uncertainties, thereby making it comparable to RMPC. In their study Drgoňa et al. Drgoňa et al. (2022) demonstrated the capabilities of DPC by applying it to a real SISO system with an unknown nonlinear plant. Experimental results showed that the performance of the controller was on par with that of an exact EMPC alternative. However, the DPC controller could handle much more complex tasks such as tracking dynamic references with dynamic constraints. Such results are promising and indeed demonstrate the ability of DPC to control nonlinear systems with disturbed, uncertain and unknown dynamics, however, the performance and scalability of DPC to MIMO systems are yet to be established.

### v3

@@ 5 Neural Network Based Optimisation for Nonlinear Model Predictive Control
Parisini and Zoppoli (1995) 1a First paper to propose the idea of training an NN to imitate MPC via pure regression. None 6 3 2 30 30 Sim
Ortega and Camacho (1996) 1c Train a shallow NN to imitate MPC using RMSE loss. Inputs to the network include past inputs and sensor readings that are not fed to MPC. None 5 3 2 7 6 Robot
Cavagnari et al. (1999) 1e Train shallow NN with sigmoid activation to imitate LMPC, generate training data from LQR solution as well as MPC. None 4 4 1 10 10 Robot
Lucia et al. (2020) 1a Train $tanh$ DNN with MSE loss to imitate MPC, then implement the approximate controller using FPGA for much faster execution. None 2 2 2 10 10 500,000x HIL
Chan et al. (2021) 1b Train DNN with MSE loss to imitate MPC, then add steady state error correction term to loss function and retrain to generate offset free DNN. None 2 1 1 Sim
Vaupel et al. (2019) 1c Train DNN to imitate MPC using MSE loss, use measurable disturbance variables as network inputs. None 7 7 1 10 8 10x Sim
@@
Nurbayeva et al. (2022) 1c MPC motion planner approximated by DNN via pure regression, explores input dimensionality reduction with autoencoders. None 6 6 6 10 10 Robot
Löhr et al. (2019) 1a Train DNN to imitate MPC with binary control inputs using RMSE loss function. None 3 3 3 24 24 12.12x Sim
Karg and Lucia (2018) 1c Train DNN to imitate MPC with binary control inputs. Include trajectories of disturbances over horizon as part of the training data. None 4 4 4 24 24 Sim
Lovelett et al. (2020) 1c Use manifold learning to map augmented state space to optimal control inputs. Then train DNN to retrieve optimal control inputs from manifold space. None 2 1 1 20 20 36,000,000x Sim
Åkesson and Toivonen (2006) 1d Train a shallow $tanh$ NN with an MPC loss function on data corresponding to closed loop trajectories. Use result to replace conventional optimiser. None 4 2 2 60 25 Sim
Drgoňa et al. (2022) 1d Create and train DNN plant based on state space architecture. Train DNN controller with loss function inspired on MPC cost. None 1 1 32 32 Robot
Li et al. (2022) 2b Formulate training data generation and DNN approximation training as a single optimisation problem. Probabilistic constraint satisfaction. 4 2 2 20 20 1,000x Sim
Parisini et al. (1998) 2c First formalised framework for generating shallow NN approximations of NMPC. Authors propose methods to compute NN size and terminal NMPC costs that make NN feasible. Bound output error and stability to region near origin. 2 2 1 100 100 Sim
Pin et al. (2010) 2c Further investigation of approximation properties of the framework proposed in (Parisini et al., 1998). Practical stability and input constraint satisfaction. 2 2 1 8 8 Sim
Pin et al. (2013) 2c Extend (Pin et al., 2010) with requirements that approximations have to meet to guarantee stability and constraint satisfaction. Practical stability and constraint satisfaction on inputs and states. 2 2 1 8 8 Sim
@@
Ahmed and Al-Dajani (1998) 1d Create a first order differentiable, MPC-inspired loss function and use it train a DNN that imitates MPC. Derivatives of loss are computed using block partial derivatives. Input constraint satisfaction via saturation and stability at equilibrium point. 2 1 1 300 300 Sim
Löhr et al. (2020) 3a Train DNN to predict first value of a binary control input and thereby enable conversion of MIQP into QP. Constraint satisfaction and input feasibility via solution of QP at every iteration. 3 3 5 48 48 3x Sim
Vaupel et al. (2020) 3a Train DNN with MSE loss to imitate MPC then use it to hot start full NMPC solution or alongside a post processing QP that corrects the outputs of the network and makes them constraint compliant. All MPC guarantees when DNN is used to hot start solver. Good practical constraint satisfaction when post processing QP is used. 6 5 3 7 5 24x Sim
Bonzanini et al. (2020) 3b Train a DNN to imitate NMPC, project outputs of DNN onto RAI which in turn will ensure that states stay within RCI. The projection operation can be precomputed as the solution of an mpQP. Input and state constraint satisfaction. 2 2 2 100 100 25 - 90x Robot
Wang et al. (2022) 1f Integrate Lagrange duality into DNN loss function to guarantee constraint satisfaction of the primal problem, train the DNN approximation through imitation learning on the modified loss function. Input and state constraint satisfaction. 3 3 2 50 50 6 - 72x Sim
@@ 5.1 Methods without guarantees
Parisini and Zoppoli (1995) were the first to propose the idea of training an NN to imitate the outputs of an NMPC controller. Their framework consisted of fitting a shallow NN with $tanh$ activation to training data generated by repeatedly applying a stabilizing NMPC controller to initial states sampled randomly from a bound set. The inputs to the network were the current system states and the outputs were the control actions calculated by the NMPC controller for the first step in the control horizon, an MSE loss function and a gradient algorithm were used to solve the regression problem. The effectiveness of the approach was demonstrated via simulation on two example problems one of which included a MIMO plant of medium size as reported in Table 2. Ortega and Camacho (1996) extended these results by applying the same method to a path tracking NMPC controller used as the navigation algorithm of a mobile robot. In this case the shallow NN was provided with additional inputs such as a parametric representation of the reference trajectory, distance sensor readings for obstacle detection, and most notably, the control actions at the previous time step. Ortega and Camacho (1996) validated their approximate controller with experiments on the real robot, thus showing, for the first time, that NNBO can indeed facilitate the deployment of NMPC controllers that would otherwise be unable to execute in real time. Cavagnari et al. (1999) further proved this point with their experimental results on a seesaw mechanism. Using the same random sampling and MSE loss framework, they replaced a stabilizing NMPC controller with an NN approximation capable of running in real time. Moreover, they introduced a small novelty by training the NN with data generated from two different controllers operating in different regimes. For samples near the equilibrium position, they used training data generated from an LQR controller with a linearised plant, for points outside this region, training data was generated using the NMPC controller. The result was an approximate NN controller that could seamlessly transition between two control behaviours for better stability. The three aforementioned studies were conducted in the late 90s. However, following their publication, the simple imitation learning approach received little attention for almost two decades.
@@
Interest in the simple imitation learning approach reignited around 2018, with many researchers proposing modifications and extensions of the method. Lucia et al. (2020) showed that the $tanh$ DNN approximation of an NMPC controller could be implemented in Field Programmable Gate Array (FPGA) hardware to achieve much faster execution of the approximate controller and thus enable the control of systems with ultra fast dynamics. The effectiveness of this approach was demonstrated on a resonant power converter application including Hardware In the Loop (HIL) testing. From these tests, Lucia et al. (2020) reported a speed up factor of approximately half a million times for their most optimised FPGA implementation compared to the original NMPC controller.
@@
Chan et al. (2021) proposed an extension of the basic framework to eliminate steady state setpoint tracking error due to approximation error. In their framework, a DNN approximation of an NMPC controller is first generated by solving the usual regression problem with an MSE loss function. Then, the approximation and original controller are applied to setpoint tracking trajectories to generate a new data set. This dataset is used to retrain the network using a modified loss function that includes an approximation error term. Retraining on this modified loss function effectively minimises the approximation error which in turn minimises, or effectively removes, setpoint tracking error. Chan et al. (2021) also implemented their approximate controller on FPGA hardware and demonstrated the effectiveness of their approach via simulation for a small problem. Alternatively, Vaupel et al. (2019) showed that acceptable setpoint tracking performance can also be achieved by including disturbance measurements in the inputs of the neural network approximation and training with the pure regression setup on an MSE loss function. For applications where disturbances can be measured, this approach is significantly simpler compared to the method proposed by Chan et al. (2021).
@@
For complex systems with difficult to parameterise inputs, Nurbayeva et al. (2022) proposed the use of autoencoders for input dimensionality reduction. In their work, an NMPC motion planner for human-robot interaction tasks was replaced with a DNN approximation and successfully tested in simulation and on the real robot. Autoencoders were used to parameterise human motion inputs to the approximate controller, robot state inputs were fed directly. While the approach may be worth considering for some applications, the authors themselves point out that the resulting controller is not generalisable to motion scenarios beyond the specific interaction for which it was trained. The controller is also unable to cope with significant deviations from the interaction it was trained for.
@@
Löhr et al. (2019) and Karg and Lucia (2018) showed that the basic imitation learning setup is also applicable to NMPC controllers with binary variables where the underlying optimisation problem is of the mixed-integer type. Löhr et al. (2019) trained their approximation using an RMSE-based loss and the inputs to their network were the current states, references and disturbances of the system. Karg and Lucia (2018) considered the same inputs but used the entire predicted disturbance trajectory over the prediction horizon instead of only the latest measurement. In both studies the effectiveness of the approach was demonstrated via simulation on heating system control problems with MIMO plants of similar size. The independence and success of these studies suggests that NNBO is a viable method for accelerating NMPC problems with binary control variables, but there is room for in-robot validation studies as well as studies that explore the more general problem of handling integer variables instead of only binary variables.
@@
In a significantly different approach, Lovelett et al. (2020) noted that the mapping function between states and optimal control actions (e.g. the function that the NN in previous approaches is trying to learn) can be complicated and difficult to learn because the augmented state space, which consists of plant states and controller references, is often a poor parametrisation of the control law. Motivated by this observation, they proposed the use of diffusion maps, a type of manifold learning algorithm, to learn a potentially lower dimensional parametrisation of the mapping function. If such a parametrisation exists, it is a simpler map between manifold space and control actions. However, using this map to obtain control actions given the augmented state is not straightforward. The authors addressed this challenge with two NNs. The first NN was used to learn the mapping between augmented state space and manifold space, and the second network learnt the mapping between manifold space and optimal control actions. In theory these regression problems are easier than directly approximating the original controller. The authors demonstrated the effectiveness of the approach via a simulation case study on a nonlinear system with a small plant. Reports on execution time improvements indicated that the neural approximation was 36,000,000 times faster than the original NMPC.
@@
So far, we have discussed methods that make use of a simple imitation learning setup where error information is used to solve a regression problem via gradient descent, hereafter we present other alternatives. Åkesson and Toivonen (2006) proposed an approach wherein the error-based loss function is replaced with an MPC-like loss function. This new loss function encourages the network to learn the dynamics of the controller instead of over-fitting it for a bound set of inputs. In their study the authors formulated this method for trajectory tracking NMPC. The network inputs are those typically available to an NPMC controller such as current estimated states, measured outputs, reference outputs, and previous control actions. The network outputs are the control command increments to be applied at the current control interval. In this case training consists of minimising the MPC-like loss function over a large set of trajectories on which the network is applied in a closed loop setup. The minimisation task can be formulated as a nonlinear least squares problem and solved using a gradient algorithm. The main advantage of this method is that the size of the resulting optimisation problem is independent of the horizon length used in the MPC loss function. This means that the NN approximation can be trained with longer horizons, surpassing what would be practically feasible for deployment in the exact NMPC controller.
@@
Drgoňa et al. (2022) built further on the idea of departing from error-based loss functions in their Differentiable Predictive Control (DPC) framework where both the optimiser and plant model of the original controller are replaced with DNNs with architectures and loss functions inspired by MPC. More specifically, the plant model is approximated using a neural state space architecture which can learn the dynamics of nonlinear systems from recorded data. The main advantage of this approach is that the model directly learns the disturbed and uncertain dynamics of the system without additional effort. The trained plant can then be used to train a policy network using an MPC-inspired loss function and a dataset that emulates closed loop control. Crucially, the policy and plant networks can be combined into a single model that is continuously differentiable making it easy to train with any modern training algorithm such as ADAM Kingma and Ba (2014). Again, the advantage of combining the policy and data-derived plant networks into a single model is that the resulting policy inherently compensates for disturbances and uncertainties, thereby making it comparable to RMPC. In their study Drgoňa et al. (2022) demonstrated the capabilities of DPC by applying it to a real SISO system with an unknown nonlinear plant. Experimental results showed that the performance of the controller was on par with that of an exact EMPC alternative. However, the DPC controller could handle much more complex tasks such as tracking dynamic references with dynamic constraints. Such results are promising and indeed demonstrate the ability of DPC to control nonlinear systems with disturbed, uncertain and unknown dynamics, however, the performance and scalability of DPC to MIMO systems are yet to be established.